A total of 50 juniors and seniors were given a mathematics test. The 35 juniors attained an average score of 80 while the 15 seniors attained an average of 70. What was the average score for all 50 students who took the test?
Answer by w_parminder(53)

Total candidates = 50
average of 35 (juniors) candidates = 80
So the sum of the marks attained by juniors = 80*35 = 2800
average of 15 (senior) candidates = 70
the sum of the marks attained by seniors = 70*15 = 1050
Total marks attained by all the candidates (seniors & juniors) = 2800 + 1050
= 3850

average score for all 50 candidates = 3850/50 = 77

That's it.....